Day #34 – 2/5/15 – Phoenix Rising

Lab Opening time: 1:30pm

At 3:00pm, the team held a meeting to make a few announcements. Today the engineering lab got new CO2 fire extinguishers that are more kind to equipment than a dry-chemical type extinguisher. We also discussed the snack room and how we have been neglecting to clean up the snack room and as a result, the snack room was off limits today. Each sub-group also got their assignment for the end of the year which is to create a document stating what each sub-group does.

Today was a busy work day with most of the team helping to assemble all of the parts we have gotten back from powder-coating and anodizing. Cynthia continued the assembling go the swerve modules and Mika, Devon, Alexander Luke, and Jack worked on the wiring of the robot.Today they mounted the Victor plate, and crimped the Anderson connectors onto the victors, they also terminated and crimped the wires from the PDP.

Top view of Victor plate
Wires from the Victors terminated and crimped
Jack and Devon working on wiring.
Jack and Devon working on wiring.
Cynthia working on the swerve assembly.
Cynthia working on the swerve assembly.

Joshua and Spian have been working on the CAD model of the the autocanners and have the designed mostly finalized. Shay continued to work on the CAD design of the elevator.

Olivia, Claire, and Mr. Widholm finished the right elevator forks today. The part take about 53 minutes to machine not including the time it took to replace the bolts and refill the coolant tank. Bahvin worked on getting the encoder mounters cut on the router.

The right elevator forks.
The right elevator forks.

The programing team worked on fixing the logger for the encoders and also worked on programing the elevator forks so that they go up incrementally instead by the height of one tote.

Business team went out to see Bob Smith Toyota about their sponsorship but were unable to meet with Mike Smith today. They also wrote thank you letters and worked on the invoices for the Ventura Regional trip. They also got back their revisions from the Clark teachers who read over their Woody Flowers Finalist Award nomination essays.

The media team worked finalizing the chairman’s award script. Chris continued the t-shirt design which needs to be done by tomorrow. The web team worked on updating the calendar on the website and are getting ready for their meeting with the drive team tomorrow.

Mr. Black and Mr. Hoard worked on continuing to reassemble the laser cutter and clean it with a 50/50 mix of alcohol and DI water, which proved to be very effective.  Lenses and mirrors were cleaned with a high-purity acetone and Kimwipes.  We are happy to report, that initial tests seem to indicate that the laser cutter has survived (!) the exothermic event (and chemical deluge) of two days prior, and appears to be functioning nominally.  So far, the laser appears to be engraving and cutting at a quality equal to if not better than before the mishap.  It certainly is now cleaner than ever before.  We still need to go through and perform a full maintenance procedure of the linear motion system, but it is fully functional currently.  A very very lucky save, and a nice way to end the day as we prepare to enter the long weekend.

Lab Closing time: 8:30pm